The greater of two numbers is 1 more than three times the smaller. Four times the greater exceeds six times the smaller by 10. Find the numbers
Answer by lwsshak3(11628)

The greater of two numbers is 1 more than three times the smaller. Four times the greater exceeds six times the smaller by 10. Find the numbers
***
let x=smaller of two numbers
3x+1=greater of two numbers
4(3x+1)-6x=10
12x+4-6x=10
6x=6
x=1
3x+1=4
smaller of two numbers=1
greater of two numbers=4